OpenStudy (anonymous):

Please help me with this question, Its the last one and i don't understand what i have to do Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data.

OpenStudy (anonymous):

Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using

OpenStudy (anonymous):

a) The coordinates of the pairs are (12, 36) and (18, 54). Slope (Gradient) = (54 - 36) / (18 - 12) = 18 / 6 = 3 It represents that for every minute the person walks, he/she loses 3 calories. b) Using (12 , 36): y = 36 + 3(x - 12) = 36 + 3x - 36 = 3x c) x = 35 minutes y = 3(35) = 105 calories

OpenStudy (anonymous):

Thank you very much!!!!!:), but how did you did it im so confused and I want to learn.. Any main tips??

OpenStudy (anonymous):

On what topic?

OpenStudy (anonymous):

On linear graphs, Remember the formulas, y = y1 + m(x - x1) where m is the slope or gradient of the line and (x1 , y1) is a point on the line. y = mx + c where m is the slope or gradient of the line and c is the y-intercept, which means the y-coordinate where the line intersects the y-axis. (y2 - y1) / (x2 - x1) = m, which is the slope or gradient, and (x1 , y1), (x2 ,y2) are 2 different points on the line.

OpenStudy (anonymous):

Thank you very much!!!!!!!!